OpenStudy (anonymous):

what is the answer to 4x-6/-2=15

OpenStudy (compassionate):

I won't give you the answers, but I can tell you how to solve it. \[\frac{4x - 6}{-2} = 5 \] You can take the inverse, and start by multiplying by -2.

OpenStudy (igreen):

\(\dfrac{4x-6}{-2} = 15\) Multiply -2 to both sides: \(4x-6 = -30\) Add 6 to both sides: \(4x = -24\) Divide 4 to both sides. What's -24 / 4? @mstevens

OpenStudy (compassionate):

15. Not 5.

OpenStudy (igreen):

It's \(\dfrac{4x-6}{-2} = 15\) not \(\dfrac{4x-6}{-2} = 5\).. @Compassionate xD

OpenStudy (igreen):

Dividing -24 / 4 will give you your final answer @mstevens

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'.

OpenStudy (anonymous):

But its not showing me the answer hell

OpenStudy (anonymous):

I like to to damn curse

OpenStudy (compassionate):

We don't give out answers here. We will guide you to an answer. Getting direct answers violates Academic Integrity. Basically, it means you're a cheater. So don't cheat. We've both given explanations. Is there anything else you need? Do you need a more direct explanation? iGreen basically broke it down to simple division for you.